Please help, I don't know how to solve this question:

A one‐year discount bond has a payout of $21,000.   If the price rises from$19,000 to $20,000, the yield to maturity ___ from ___ % to ___ %?

Kindly show me the step please. Thank you.
Answer #1

Yield to maturity = (payout - price)/ price × 100

Yield to maturity when the price is $19000 = (21000 - 19000) / 19000 × 100 = 10.52%

Yield to maturity when the price is $20000 = (21000 - 20000)/20000 × 100 = 5%

So, the yield to maturity decreases from 10.52% to 5%.